OpenStudy (anonymous):

Show an equation and solution for the problem. Jasmine and Keyla live in the same town. At 8:00 a.m. Jasmine drives west at 60 mph and two hours later Keyla drives east averaging 65 mph. How many hours after Jasmine leaves will they be 650 miles apart?

OpenStudy (anonymous):

(60(x+2))+65x=650 and solve for x 60x+120+65x=650 125x+120=650 125x=530 x=4.24 4 hours and 14.4 minutes